In Desmos:

Describe the transformations applied to the parent f(x)= 3sqaure root of x into the new equation f(x)=-3sqaure root of x into the new equation f(x) = - 3sqaure root of x-5-3

1
Expert's answer
2021-03-09T06:01:27-0500

Solution.

"y=\sqrt[3]{x}" (blue line)

"y=-\sqrt[3]{x}" - symmetry about the x-axis (green line)

"y=-\sqrt[3]{x-5}" - parallel transfer by 5 units to the right (purple line)

"y=-\sqrt[3]{x-5}-3" - parallel transfer by 3 units down (red line)